OpenStudy (anonymous):

[4.07]What is the slope of the line passing through the point (3, −4) and perpendicular to the line y = 3x + 5? Write your answer as a fraction

OpenStudy (anonymous):

Do you need the slope only? That will be -1/3. If you need the equation, substitute the point (3,-4) into y=-1/3x + c to get c

OpenStudy (anonymous):

yes please

hartnn (hartnn):

u get -1/3 because the product of slopes of perpendicular lines is -1 so x*3=-1 so x=-1/3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you help me with one more?

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

[7.03] The length of each side of a square wooden box, in inches, is represented by the expression 5z4. The volume of the box, in cubic inches, is (5z4)3. Which simplified expression represents the volume of the box? 125z64 125z12 125z7 15z12

hartnn (hartnn):

 whats cube of 5?